### v1

# A Heuristic Philosophical Discourse on Various Applications of

Abstract Differential Geometry in Quantum Gravity Researchthanks: This paper is a philosophical distillation of the basic concepts and central results of this author’s ongoing research project, spanning the last three decades, of applying Anastasios Mallios’s Abstract Differential Geometry to the ‘persistently stubborn’ problem of formulating a conceptually sound, mathematically consistent and calculationally finite Quantum Theory of Gravity. The paper will be posted at the General Relativity and Quantum Cosmology website www.arXiv.org/gr-qc before October 2024. This work may be viewed as the philosophical résumé and aftermath (:after the Maths!) of the following published papers [62, 63, 64, 73, 74, 75, 76, 77, 78, 79, 84, 85], as well as a sequel to the Dodecalogue paper [79] in the aftermath of the recent publication [80] and the pre-prints [81, 82, 83], which are currently work-in-progress in the pipeline. In turn, a longer version of the paper will constitute a chapter in a research monograph type of book that we had and have been working on, in collaboration with the late Professor Anastasios Mallios, since 2003 [65].
Ioannis Raptis Supply & Substitute Secondary School Teacher of Mathematics, Physics and Chemistry, Reeson Education, London, United Kingdom; email: irapti11@gmail.com
(Thursday, 15th of August 2024
(Version I))
Abstract
In the present paper, we outline and expound the fundamental and novel qualitative-cum-philosophical premises, principles, ideas, concepts, constructions and results that originate from our ongoing research project of applying the conceptual panoply and the technical machinery of Abstract Differential Geometry (ADG) to various persistently outstanding issues in Quantum Gravity (QG) [62, 63, 64, 73, 74, 75, 76, 77, 78, 79, 84, 85]. This paper may be regarded as a sequel to the paper [79] in the aftermath of the paper [80], which is currently in press. At the end of the paper, we discuss the potential philosophical repercussions of two possible future research routes that the main stream of our applications of ADG to QG may bifurcate towards in view of three independent, but overlapping, research papers that are currently under development [81, 82, 83].
PACS numbers: 04.60.-m, 04.20.Gz, 04.20.-q
Key words: Natural Philosophy of Quantum Gravity and Quantum Yang-Mills Gauge Theories; Abstract Differential Geometry; sheaf theory; sheaf cohomology; category theory; topos theory; geometric prequantisation; canonical quantisation; background spacetime manifold independence; quantum gravity as a purely quantum gauge theory; gauge theory of the third kind; third sheaf cohomological quantisation of gravity and gauge theories
1 Prolegomena-cum-Motivation: Why Adopt a Philosophical Stance in Quantum Gravity Research?
Quantum Gravity (QG), very broadly speaking, is an attempt to unite the laws of Physics that describe dynamics at large (:cosmological) scales—as encoded in Einstein’s General Relativity (GR) equations for the gravitational field that guides the motion of large material objects, with the dynamical laws for the other three fundamental forces that guide matter fields and their quantum particles at small (:subatomic) scales—as encoded in Quantum Theory and its application to Special Relativistic field physics, commonly known as Quantum Field Theory (QFT). [^1] In this paper, we use the terms QFT and Quantum Gauge (QGT) or Quantum Yang-Mills Theories (QYMT) of matter interchangeably.
There is currently a plethora of various and glaringly diverse approaches to QG, and it is not the ‘vain’ or ‘quixotic’ aim of this paper to list them all herein, let alone to review their conceptual and technical import or their successes and shortcomings. On the one hand, there is no unanimous agreement on what the ‘right’ or ‘correct’ approach to QG is and the diversity of the various different approaches—from string theory [25, 91], [^2] For current developments in this field, see the second reference above [91]. to loop quantum gravity [2, 3, 4, 44] [^3] For an exhaustive modern exposition of loop quantum gravity, see the last reference [44] above. and causal sets [6, 93, 94, 95], for example—exemplifies exactly that.
On the other hand, it would be informative for the reader to list a triplet of very general predicates and characteristics—theory making ‘imperatives’ as it were—that the desired and hitherto elusive Quantum Theory of Gravity should possess, properties that all the aforementioned diverse approaches to QG aim or aspire to satisfy in one way or another:
  1. 1.
    Economy (E): Conceptual clarity and depth, as well as economy and simplicity of the underlying fundamental physical theory construction principles;
  2. 2.
    Mathematical Consistency (M): Internal mathematical consistency (:self-consistency), mathematical representability and predictive power, technical innovation and efficacy, wide ranging utility and versatility of application, and, as a bonus, abstract mathematical simplicity and beauty; [^4] With respect to this last point, we tacitly assume that‘beauty, especially of the mathematical kind, is in the eye of the beholder’—in this case, of the theoretician/mathematician—hence it is largely of a subjective aesthetic nature.
  3. 3.
    Finiteness (F): Calculational freedom from unphysical (observable) infinities, anomalies and ‘singularities’ of all kinds at which the aforementioned physical laws might seem to break down, hence are deemed to become unphysical, uninformative, unpredictive and, ultimately, practically useless and therefore obsolete.
We will henceforth refer to the triplet of desirable QG traits above by the acronym EMF1.
‘Mission Statement’: Our philosophical discourse in the present paper will focus on showing, arguing and discussing, with extensive references to the existing published literature, that the ADG-theoretic approach to QG, hitherto to be referred to as ADG-gravity, goes a long way towards satisfying the EMF1 triptych of theory making imperatives above.
At the same time, the basic ‘justification’ for engaging in a philosophical discourse about the import of ADG-gravity in QG research is another triplet EMF2:
  1. 1.
    Explication (E): Explanation and interpretation of new concepts, techniques and results from applying ADG in QG research;
  2. 2.
    Mathematical Efficacy (M): Discussion of the mathematical power of ADG in addressing and resolving certain key QG problems and issues associated with EMF1;
  3. 3.
    Future Prospects and Developments (F): Discussion of future prospects for QG theory growth and development, as well as anticipation in what direction will QG research move in view of the new ideas and theoretical paradigms that ADG brings forth.
Expository Declaration 1: In the sections that follow, whenever we discuss and analyse a qualitative-philosophical aspect of ADG-gravity, at the end of the discussion we will mark it by boldface markers in brackets like, for example, (F1): this would mean that the qualitative/philosophical characteristic of ADG-gravity being analysed and discussed satisfies the Finiteness 1 aspect of the EMF1 triplet above.
We would like to kick-off our philosophical discourse on applications of ADG in QG with a very telling quote of Gerard ’t Hooft just after the turn of the new millennium [98], which kind of gives a raison d’être and a raison de faire to our endeavours herein:
(Q1)                     …The problems of quantum gravity are much more than purely technical ones. They touch upon very essential philosophical issues [^5] Our emphasis.
the basic idea behind the quotation above is that it motivates us to go beyond the ‘technicalities’ of various formal conceptual and mathematical issues in current QG research, and discuss the deeper semantics and philosophical nature underlying, or even possibly transcending, those conceptual and mathematical ‘technicalities’. Which brings us to the second Expository Declaration of this paper.
Expository Declaration 2: It is a conscious decision and choice of this author not to include a single quantitative expression (e.g., technical mathematical formula) in the present paper, which is purely of a qualitative (discursive) character. For concise definitions of formal technical concepts and their associated mathematical formulae/modelling/equations, the reader will be directly referred to the relevant published literature.
Expository Declaration 3: Our philosophical treatise below is organised in short sections that end with ‘Aphorisms’—short ‘statements’ that distill the main philosophical gist of each section.
2 General Relativity is formulated by the Classical Differential Geometry on a Pointed Smooth Base Spacetime Manifold
$\bullet$ General Relativity (GR), the classical theory of gravity, is inextricably tied to a $\mathcal{C}^{\infty}$-smooth (alias, differential) base spacetime manifold $M$ for its mathematical formulation via Classical Differential Geometry (CDG)—the Newtonian Calculus based geometry of differential manifolds.
$\bullet$ In turn, the differential manifold $M$, as a geometrical point-set, is equivalent to the structure algebra sheaf $\mathcal{C}^{\infty}(M)$ of germs of smooth coordinate functions of its points.
$\bullet$ In GR, the dynamical law of gravity is formulated as Einstein’s nonlinear partial differential equations of a $\mathcal{C}^{\infty}$-smooth spacetime metric $g_{\mu\nu}$ (and its derivatives), whose $10$ components are supposed to represent the gravitational field potentials.
$\bullet$ Accordingly, GR’s Priciple of General Covariance (PGC) is mathematically represented by the group $\mathrm{Diff}(M)$ of (active) diffeomorphisms of $M$ [^6] By definition, a diffeomorphism of a smooth manifold $M$ is an automorphism of $M$ that preserves its differential geometric (:smooth) structure, as the latter is effectively encoded in the structure algebra sheaf $\mathcal{C}^{\infty}(M)$ of smooth coordinate functions of $M$’s point events, as noted above.
Aphorism 1. GR uses CDG to formulate the gravitational dynamics as a differential equation for the smooth metric (and its derivatives) on a background geometrical differential spacetime manifold $M$. In turn, the PGC of GR is represented by the spacetime diffeomorphism group $\mathrm{Diff}(M)$ of the underlying $\mathcal{C}^{\infty}$-smooth manifold $M$.
2.1 Feynman’s ‘Fancy-Schmanzy’ Differential Geometry and Isham’s No-Go of Differential Geometry in Quantum Gravity
Arguably, the smooth background geometrical spacetime manifold, whether curved or flat, is responsible for both the singularities of the smooth gravitational field $g_{\mu\nu}$ of GR [22, 23, 27, 28, 29, 30, 71], as well as for the pestilential non-renormalisable unphysical infinities that assail QFT on flat Minkowski spacetime [26, 5].(F1)
Mainly due to these pathologies, QG researchers as early as Feynman and as recently as Isham, have questioned altogether the use of CDG on a smooth spacetime manifold, whether flat (QFT) or curved (GR), as the appropriate mathematical framework via which to formulate QG.
In this line of thought, we first recall from Bryan Hatfield’s Quantum Gravity Foreword to Feynman’s Lectures on Gravitation [19] the following telling excerpt:
(Q2)                     “…Thus it is no surprise that Feynman would recreate general relativity from a non-geometrical viewpoint. The practical side of this approach is that one does not have to learn some ‘fancy-schmanzy’ (as he liked to call it) differential geometry in order to study gravitational physics. (Instead, one would just have to learn some quantum field theory.) However, when the ultimate goal is to quantize gravity, Feynman felt that the geometrical interpretation just stood in the way. From the field theoretic viewpoint, one could avoid actually defining—up front—the physical meaning of quantum geometry, fluctuating topology, space-time foam, etc., and instead look for the geometrical meaning after quantization… Feynman certainly felt that the geometrical interpretation is marvellous, ‘but the fact that a massless spin-$2$ field can be interpreted as a metric was simply a coincidence that might be understood as representing some kind of gauge invariance’ [^7] Our emphasis of Feynman’s words as quoted by Bryan Hatfield.…”
And he further added categorimatically in [18] that:
(Q3)                     …the simple ideas of [differential] geometry, extended down to infinitely small, are wrong.
While more recently, Chris Isham firmly posited in [31]:
(Q4)                     …at the Planck-length scale, differential geometry is simply incompatible with quantum theory…[so that] one will not be able to use differential geometry in the true quantum-gravity theory…
We may distill the above to our second Aphorism:
Aphorism 2. All the anomalies and pathologies of GR and QFT in the form of singularities and other unphysical infinities originate from the a priori assumption of a background differential manifold as a geometrical model for spacetime.(F1)
Thus, we can combine Aphorisms 1 and 2 to the following ‘vicious circle’ statement:
Fundamental Vicious Circle. If we wish to formulate the dynamical laws of QG as differential equations proper, it seems that we have to use the concepts and techniques of CDG on a smooth manifold. However, the latter is responsible for both the singularities of GR and the unphysical the infinities of QFT—sites in the spacetime manifold where the laws of physics appear to break down or lead to unphysical infinities for important observable quantities; hence, we seem to arrive at an impasse. [^8] The reader should note here that, in the three quotes above, both Feynman and Isham question the Mathematical Efficacy of CDG in QG (M1,F1, M2).
3 Enter ADG
Below, we itemise the basic tenets of ADG:
$\bullet$ Abstract Differential Geometry is more of a Leibnizian (:relational), rather than Newtonian (:geometrical), purely algebraic (:sheaf-theoretic and homological algebraic) way of doing differential geometry (:Calculus) [15, 8], without at all recourse to or dependence on a pointed, smooth geometric locally Euclidean background space (:a $\mathcal{C}^{\infty}$-smooth manifold) for its concepts, technical machinery and constructions thereof [47, 48, 53].(M2)
$\bullet$ The most fundamental concept of ADG is that of connection (viz. generalised differential) $\mathcal{D}$ acting on a vector sheaf $\mathbf{\mathcal{E}}$ over a suitably algebraized, by a certain so-called algebra structure sheaf $\mathbf{A}$ of generalised coordinates or what Mallios coined ‘arithmetics, arbitrary topological space $X$ as a linear and Leibnizian sheaf morphism. The pair $(\mathcal{D},\mathbf{\mathcal{E}})$ is coined an ADG-field.(M1,M2)
$\bullet$ Based on the concept of connection, ADG erects the whole edifice of CDG (plus more), but in the manifest absence of a background geometrical manifold.
$\bullet$ Thus, a new, entirely algebraic (:relational) notion of geometry emerges, whereby, geometry does not pertain to the configuration states (‘shape’) and measurement of objects living in an a priori posited (:postulated) ether-like background space [20], but rather, it derives from the algebraic (:dynamical) relations between the objects that live on that ‘space’. [^9] These ‘objects’ are the very ADG-connection fields acting on the sections of the vector sheaves involved.
$\bullet$ The last bullet point put in more physical terms: physical geometry (:physical ‘spacetime’) is not a priori posited like the differential spacetime manifold $M$ of GR. Rather, it derives from the dynamics (:the differential equations involving the connection field $\mathcal{D}$) of the ‘objects’ (:the dynamical physical fields themselves).(E2,M2)
We may distill the essential gist of the bullets above to the following Aphorism:
Aphorism 3. From the ADG-theoretic perspective, physical geometry (or physical ‘spacetime’) derives from, or is the outcome of, the algebraic dynamical relations between the ADG-fields (:the physical laws, which are formulated categorically as equations between he relevant sheaf morphisms that the ADG-connection fields correspond to). [^10] In this regard, one may think of the more commonly used mathematical term ‘solution space’ derived from a set of (differential) equations set up (on a manifold) by the usual CDG-means. That ‘solution space’ is the ‘physical geometry’.(E2,M2)
4 The Point of Pointlessness and Finiteness: the ADG Evasion of Spacetime Singularities and the Management of Infinities
$\bullet$ In the homological algebraic (:category-theoretic) setting of ADG, the singular, ideal and physically unrealistic notion of a geometrical point [^11] The ‘physically unrealistic’ nature of a geometrical point (:an ideal spacetime event, so to speak) can be appreciated if one considers the fact that one cannot localise an ‘event’ (:measure the value of, say, the gravitational field at a point) with more accuracy than the Planck length without creating a black hole (:think for example of the inner Schwartzschild singularity right at the point-mass source, where the gravitational field blows up without bound. is meaningless.(F1)
$\bullet$ Mutatis mutandis for the continuous infinity of point-events that the smooth spacetime manifold accommodates: for example, the non-renormalisable infinities of QFT in Minkowski spacetime effectively arise from the fact that one can in principle pack an uncountable infinity of events (:field values) in a finite spacetime volume.(F1)
$\bullet$ In both the point of pointlessness and finiteness, ADG has been applied towards formulating on the one hand a locally finite, causal and quantal version of Lorentzian vacuum Einstein gravity and free Yang-Mills theories, and on the other, the same dynamical equations are seen to hold over spaces that are everywhere dense with singularities of the most unmanageable kind from the point of view of ADG [50, 51, 55, 56, 62, 63, 64, 66, 67, 68, 75].
$\bullet$ The results above are due to the purely algebraic and background pointed continuous spacetime manifold independent character of ADG [47, 48, 53, 79].
Again, we may distill the essential gist of the bullets above to the following Aphorism, our fourth one:
Aphorism 4. We can formulate the dynamical equations of Einstein and Yang-Mills over highly pathological and problematic spaces, especially when viewed from the smooth background spacetime manifold perspective of CDG. Thus, singularities (and their associated infinities) are not insuperable obstacles and ‘sites’ where the differential equations that represent the dynamical field laws of Nature appear to break down. Not only we can evade them by ADG-theoretic means, but also we can ‘calculate’ (:do Calculus!) in their very presence, in spite of them. The inherently algebraic differential geometric mechanism of ADG is genuinely background smooth spacetime independent, hence it does stumble on its inherent anomalies and pathologies [50, 51, 55, 56, 62, 63, 64, 66, 67, 68, 75]. (E1,M1,F1,E2,M2)
Thus, by ADG-theoretic means we are able to evade the ‘vicious circle’ statement that we made earlier after the first two aphorisms, as well as to question both Feynman’s (Q2,3) and Isham’s (Q4) doubts about using differential geometric ideas in QG research.
4.1 ADG Gravity: Einstein’s Purely Algebraic Description of Reality in the Quantum Deep
In the Philosophy of Physics there is a well established view that Einstein in vain pursued his Unified Field Theory [^12] Einstein had originally coined it Unitary Field Theory instead [17]. research on a continuous spacetime manifold in spite of the inherently finitistic and algebraic description of Physical Reality at subatomic scales that Quantum Theory brought about.
The following couple of quotations from Einstein’s The Meaning of Relativity [17] testify to that:
(Q5)                     …One can give good reasons why reality cannot at all be represented by a continuous field. From the quantum phenomena it appears to follow with certainty that a finite system of finite energy can be completely described by a finite set of numbers. [^13] Our emphasis. This does not seem to be in accordance with a continuum theory, and must lead to an attempt to find a purely algebraic theory for the description of reality [^14] Our emphasis.
and a similar quote from [17] that also mentions singularities:
(Q6)                     …Is it conceivable that a field theory [^15] Here, Einstein was implicitly alluding to his Unitary Field Theory, which, according to his vision, could hopefully ‘explain away’ quantum phenomena. permits one to understand the atomistic and quantum structure of reality? Almost everybody will answer this question with ‘no’. But I believe that at the present time nobody knows anything reliable about it. This is so because we cannot judge in what manner and how strongly the exclusion of singularities reduces the manifold of solutions. We do not possess any method at all to derive systematically solutions that are free of singularities [^16] Our emphasis.[17]
In connection with the two Einstein quotes (Q5,6) above, with our fifth Aphorism next, which closes this section, we kill two birds with one stone:
Aphorism 5. ADG is an entirely algebraic method for formulating gravity and quantum Yang-Mills theories of matter field theoretically and finitistically by evading singularities and without any dependence on a background geometrical spacetime continuum, with its all inherent singularities and associated unphysical smooth field infinities [64, 75, 78, 79, 80].
4.2 Sheaf Theory and the Transition from Local to Global
At the basis of ADG (alias, The Geometry of Vector Sheaves), lie the purely algebraic methods of sheaf theory [7, 47, 48, 53]. Unlike the geometry of smooth vector bundles, which features prominently in the geometrisation of Physics that gauge theory brought about [24], sheaf theory has been slow in coming in QG research.
Structures closely related to sheaves are special type of categories called topoi [46, 45], which are pointless spaces having their own internal logic. Topoi have been applied to both quantum logic [9, 10, 11, 12] and quantum spacetime structures, including QG research [35, 36, 37, 38, 39, 40, 76, 77, 80, 69, 70]
In this regard, very early on, Rudolph Haag [26] intuited the great import that sheaf theory could bring to QFT as sheaves are structures tailor-cut to encode and transmit information (from local measurements of quantum observables, for example) from local to global in QFT:
(Q7)                     Germs. We may take it as the central message of Quantum Field Theory that all information characterizing the theory is strictly local i.e. expressed in the structure of the theory in an arbitrarily small neighborhood of a point. [^17] Our emphasis. For instance in the traditional approach the theory is characterized by a Lagrangean density. Since the quantities associated with a point are very singular objects, it is advisable to consider neighborhoods. This means that instead of a fiber bundle one has to work with a sheaf. The needed information consists then of two parts: first the description of the germs, secondly the rules for joining the germs to obtain the theory in a finite region [^18] Again, emphasis is ours.
Indeed, the vector and algebra sheaves involved in ADG and their associated topoi have been used very successfully in analysing the structure of the algebras of local quantum observables and how these stitch up from local to global [105, 106]. Moreover, the ADG sheaves and their associated topoi have been applied to address important issues in QG research [76, 77, 79, 70, 69, 80].
5 Revisiting Feynman: Gravity as Gauge Theory
Returning to the Feynman quote in Section 2, we wish to dwell a bit on his remark that the fact that the gravitational field was identified with the smooth spacetime metric $g_{\mu\nu}$ of the CDG-based Riemannian geometry on a differential spacetime manifold, was an ‘accident’ of theory making. [^19] That is to say, Einstein formulated GR as the dynamics of the metric, ‘because’ he used the CDG-based Riemannian geometry of a smooth base spacetime manifold [71]. Rather, Feynman intuited that:
The deeper character of gravity is that it is a gauge force, much like the other three fundamental forces, while the methods of CDG would be ineffective in the QG deep. [^20] Similarly to what Isham said in the quote following Feynman’s.
In subsequent developments in GR, we were able to cast gravity as a gauge theory in the new Ashtekar variables involving a spin-Lorentzian gravitational connection [2] and apply the new, first-order formalism [^21] The Ashtekar formalism in terms of the tetrad $e_{\mu}$ and the spin-connection $\mathcal{A}_{\mu}$ dynamical variables is coined first order, while the GR of Einstein is based solely on the smooth spacetime metric $g_{\mu\nu}$ as its sole dynamical variable. to a new candidate for (canonical) QG called Loop Quantum Gravity [3, 44]. Albeit, on the one hand, the metric was still implicitly involved in the dynamics in the guise of the vierbein comoving tetrad $e_{\mu}$, while the smooth spacetime manifold was still present as a geometrical background in order for the canonical formalism to be applied by the methods of CDG. [^22] With all the problematic issues and pathologies that this dependence causes to QG (such as the $\mathrm{Diff}(M)$ constraint problem and the problem of time [32, 33, 34].
The words in the quote above leave us with the following quandary, posited below as a rhetorical question in the light of ADG:
Is there a way to view gravity as a gauge theory, as Feynman inuited and envisaged, while still be able to apply to it differential geometric ideas, methods and techniques in spite of Feynman’s and Isham’s No-Go of CDG in QG research?
Which brings us to the next subsection about the ADG perspective on gravity as a gauge theory.
5.1 Enter ADG: Gravity as Pure Gauge Theory of the Third Kind
One of the central results of the application of ADG to QG is that:
Aphorism 6: Gravity is a pure gauge theory, without recourse to an underlying (smooth) spacetime manifold structure for either its mathematical formulation or its physical interpretation. Gravity involves the dynamics of the ADG-theoretic Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, which is simply a connection $\mathcal{D}$ on a suitable vector sheaf $\mathbf{\mathcal{E}}$. The dynamical Einstein equations are derived from a variational principle applied on the ADG-version of the Einstein-Hilbert action functional [53, 64, 80]. [^23] By ‘pure gauge theory’ above, it is meant that the sole dynamical variable in the theory is the gravitational connection $\mathcal{D}$, acting on the sections of a suitable vector sheaf $\mathbf{\mathcal{E}}$, and nothing else. The corresponding formalism has been coined half-order formalism, to distinguish it from the first-order formalism of Ashtekar and its Loop QG outgrowth, and of course from the usual second-order formalism of the original Einstein theory (GR). (E1,M1,F1)
6 Field Solipsism and Functoriality: The Point of Spacetimelessness, Generalised Principle of General Covariance and a Different Perspective on the ‘Measurement Problem’
A geometrical point is mathematically an ideal and physically an unrealistic (:singular) entity. We discussed earlier how the sheaf-theoretic ADG and its pointless topos-theoretic extension evade the pointed background geometrical spacetime continuum of events of GR. We also noted how the ADG-gravitational connection field is the sole dynamical variable in the theory, while the underlying spacetime metric $g_{\mu\nu}$ of the usual CDG-based GR is manifestly absent from our theory. This on the one hand is supposed to depict the pure gauge character of ADG-gravity à-la Feynman, and on the other, to support the aforementioned ADG-field solipsism: that is to say, that
the ADG-gravitational dynamics does not need or depend at all on a background differential spacetime manifold for either its differential geometric formulation as a differential equation proper, or for its physical interpretation. ADG-gravity is a genuinely background independent theory. The result is that the sole dynamical variable in ADG-gravity is the gravitational Einstein ADG-field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, a feature that has been coined field solipsism and the Einstein-Hilbert variational action principle dynamics that $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ obeys has been coined ADG-field autodynamics (:autonomous gravitational field dynamics), with no dependence whatsoever on a background spacetime manifold with its inherent gravitational singularities and unphysical field infinities [64, 76, 78, 79, 80].(E1,M1,F1,E2,M2)
The discussion above brings to the forefront one very telling Einstein quote from [16]:
(Q8)                     Time and space are modes by which we [^24] Our emphasis. think, not conditions in which we live.
Space and time are human inventions convenient for representing, localising and quantifying our measurements of physical observable entities.
Which brings us to the idea of spacetime point coordinates, or equivalently, spacetime determinations/localisations/measurements of events or field-values. The locution of every point field-value or event in the spacetime manifold of GR is supposed to be determined (:measured) by four real spacetime coordinates (:coordinate functions) with respect to a given coordinate system (measurement frame of location).
The Principle of General Covariance (PGC) of GR mandates that the law of gravity (:Einstein’s equations) is generally covariant; that is to say, it is invariant under any arbitrary general coordinate transformation. [^25] Technically, we say that the group of symmetries of GR is $GL(4,\mathbb{R})$, the group of general linear transformations of the locally Euclidean (:$\mathbb{R}^{4}$) spacetime manifold $M$. Equivalently, we may state the PGC of GR as the Kleinian symmetry group of the background $\mathcal{C}^{\infty}$-smooth spacetime manifold $M$, as follows:
The symmetry group of GR is the group $\mathrm{Diff}(M)$ of differentiable automorphisms (:diffeomorphisms) of the background smooth spacetime manifold $M$.
By contrast, in ADG-gravity, where we have no background spacetime manifold and the ADG-field autodynamics is purely gauge,
The gauge symmetry group sheaf is the principal sheaf ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of structure sheaf $\mathbf{A}$-automorphisms of the associated vector sheaf $\mathbf{\mathcal{E}}$ of the ADG-Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ [64, 76, 78, 79, 80].
whereby, as noted earlier, $\mathbf{A}$ is the structure sheaf of generalised arithmetics or coordinates in the theory.
In other words, the ADG-gravitational dynamics, which is formulated entirely categorically in terms of the connection $\mathcal{D}$ sheaf morphism, is respected by (:‘remains invariant under’) all our generalised measurements (:arithmetics, event coordinate determinations) encoded in the structure sheaf $\mathbf{A}$. The gravitational dynamics ‘sees through’ all our coordinate measurements (:spacetime event localisations) in $\mathbf{A}$ [79, 80].
6.1 The Issue of Functoriality
In [80] it has been shown that the aforesaid PGC of GR, which is tantamount to $\mathbf{A}$-invariance in ADG-gravity, is an example of the Functoriality of ADG-gravity. In other words, since the dynamics is categorically represented as equations involving the connection sheaf morphism $\mathcal{D}$ of the ADG-Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, the PGC is represented by functors that preserve the relevant categories. [^26] In category-theoretic parlance, a functor between two categories, is a map or transformation that respects the objects and arrows of the two categories. In [80], in continuation and extension of [58, 59, 60], it was shown and argued that the relevant functors are, in fact, special types of $\mathbf{A}$-preserving functors (or $\mathbf{A}$-morphisms) called geometric morphisms, which preserve the ‘geometric’ structure of the vector sheaves involved in the dynamics.
We distill all the foregoing discussion into our seventh Aphorism below:
Aphorism 7. In ADG-Gravity, the dynamics is purely gauge and background spacetime manifold independent and functorial, while the PGC is functorially represented in terms of the principal group sheaf of automorphisms of the relevant vector sheaves as $\mathbf{A}$-invariance
citemall14,mall15,mall13,rap15. The latter simply means that the ADG-field dynamics, which in its Einstein-Hilbert action expression involves the curvature of the connection which is an $\otimes_{\mathbf{A}}$-tensor, remains invariant (or ‘unperturbed’) by our generalised coordinate ‘measurements’ that are organised in the structure algebra sheaf $\mathbf{A}$.(E1,M1,F1,E2,M2)
Thus, all our generalised measurements are represented in ADG as sections of the structure algebra sheaf $\mathbf{A}$ [^27] The reader should note here that the structure sheaf $\mathbf{A}$ is supposed to be a sheaf of abelian (:commutative) algebras. This reflects our primitive assumption that we always measure commutative numbers, while in the Quantum Theory it is supposed to be the ADG-version of Bohr’s Correspondence principle: although quantum observables may be noncommutative q-numbers, our measurements thereof are commutative c-numbers. In turn, the ADG-gravitational dynamics, since it is $\mathbf{A}$-functorial, [^28] Or as Mallios originally coined it: $\mathbf{A}$-invariant. is not ‘disturbed’ at all by our generalised field measurements in $\mathbf{A}$. Furthermore, the $\otimes_{\mathbf{A}}$-functorial ADG-gravitational field dynamics does not break down in any differential geometrical sense in the presence of any type of singularity that may be encoded in the structure sheaf $\mathbf{A}$.
The pair of observations above, namely that:
  1. 1.
    The ADG-gravitational dynamics is unperturbed by our generalised measurements in $\mathbf{A}$; and,
  2. 2.
    The ADG-gravitational dynamics does not break down in any (differential geometric) sense by any kind of ‘singularities’ or ‘anomalies’ present in $\mathbf{A}$,
reflect what we have elsewhere called the Principle of ADG-Field Realism [78, 79, 80]. Which brings us to the last section.
7 Gauge Field Theory of the Third Kind and its Third Quantisation
The last philosophical issue of ADG-gravity that we would like to discuss in this paper is two-fold:
$\bullet$ Gauge Field Theory of the Third Kind. We discussed earlier how from an ADG-theoretic perspective gravity is regarded as a gauge theory. We noted that the ADG-formalism may be coined half-order formalism, to distinguish it from the original second-order formalism of Einstein, whereby the dynamical variable is the smooth Riemannan spacetime metric $g_{\mu\nu}$, or from the more recent first-order formalism of Ashtekar, whereby the gravitational dynamical variables are the spin-connection $\mathcal{A}_{\mu}$ and the vierbein frame $e_{\mu}$. In ADG-gravity, the sole dynamical variable is the Einstein (connection) field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
As such, the ADG-based gauge-theoretic formulation of gravity, without recourse to any background spacetime manifold, has been called pure gauge field autodynamics [64, 78, 79, 80].
The second denomination gauge theory of the third kind comes from the observation that the first $U(1)$ gauge (or scale) theory for the electromagnetic and the gravitational field due to Weyl [103] was a global gauge theory, [^29] In the sense that Weyl showed that non-spacetime localised (global) gauge/scale invariance implies the conservation of electric charge in much the same way that general coordinate invariance leads to the conservation of energy and momentum in gravitational dynamics. while the gauge theories underlying the three fundamental forces (other than gravity) of the Standard Model [^30] That is, the electromagnetic (with local gauge group $U(1)$), the weak nuclear (with local gauge group $SU(2)$) and the strong nuclear (with local gauge group $SU(3)$) forces. are flat Minkowski space localised gauge theories [24].
By contrast, aside from its half-order formalism, our ADG-based gauge-theoretic formulation of gravity (:ADG-gravity), although local by its sheaf-theoretic character, is not background spacetime localised, since there is no background spacetime manifold to localise and solder it on to begin with.(E1,M1, F1,E2,M2)
$\bullet$ Third Quantisation. In [78, 79], and recently in [80], a third canonical type of ADG-field quantisation scenario was proposed according to which certain local, sheaf cohomological characteristic forms for both the vector sheaf part $\mathbf{\mathcal{E}}$ and the connection part $\mathcal{D}$ of the ADG-theoretic vacuum Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ were seen to obey canonical type of Heisenberg non-commutation relations, albeit, explicitly not parametrised by a background spacetime manifold. [^31] In the sense that they are not equal-time commutation relations which, in the usual canonical QG scenario, would have been required to obey some global hyperbolicity type of foliation of the background spacetime manifold into time-parametrised 3-dimensional spacelike hypersurfaces. This is to be expected as our ADG-gravity does not depend at all on an external (background) spacetime manifold, as well as to be desired, as our third quantisation scenario would be expected to ‘algebraically close’ within the autonomous and ‘solipsistic’ ADG-theoretic Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
We distill these remarks to the following eighth Aphorism:
Aphorism 8. There are no external geometrical structures, such as a background spacetime manifold, in our theory: all there is is $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ and its Yang-Mills counterparts $\mathcal{F}_{YM}=(\mathbf{\mathcal{E}},\mathcal{D})$; hence, if the autonomous (:autodynamical) ADG-fields are to be quantum (or quantised) in any way, they should be quantum (or quantised) from within themselves, not from without [78, 79, 80].(E1,M1,F1,E2,M2)
8 Brief Philosophical Musings on the Future of ADG-Gravity
This author’s current research on ADG-gravity focuses on the following three fronts:
  1. 1.
    To organise the recently discovered ‘time-asymmetric algebras’ in [81] [^32] These algebras originally appeared, in primitive form, in this author’s Ph.D. thesis [72], in which the early seeds for a time-asymmetric quantum spacetime structure and gravity were planted. into vector sheaves à-la ADG and, by employing the rich differential geometric mechanism of ADG, explore the possibility of developing a time-asymmetric Dirac equation on the resulting sheaves, possibly with ADG-gravity coupled to it [83].(F2)
  2. 2.
    The project above dovetails snugly with our current musings in [82], where we apply ADG to develop a time-asymmetric version of the vacuum Einstein equations for a finitary spin-Lorentzian gravitational connection [73, 74, 62, 63, 64] on Finkelstein’s quantum net as originally worked out by Steve Selesnick [86, 87, 88, 89, 90], and relate this asymmetry to the fundamental asymmetry that Penrose has for many years anticipated that the true QG theory should account for [41].(F2)
  3. 3.
    The main philosophical query that will arise from the three papers above [81, 82, 83] is that the anticipated fundamental time-asymmetry of the true QG theory may not only be traced back to time-asymmetric initial conditions for the Universe, [^33] Like Penrose’s Weyl curvature hypothesis. but also it may be due to the fundamentally time-asymmetric quantum gravitational dynamics themselves (:time-asymmetric vacuum Einstein equations for ADG-gravity).(F2)
The quest continues…
Acknowledgments
I am greatly indebted to Professors Goro Kato (Department of Mathematics, California Polytechnic Institute, San Luis Obispo) and Steve Selesnick (Department of Mathematics, University of St Louis, Missouri) for numerous stimulating exchanges on a plethora of topics in Mathematics, Physics, Philosophy and Poetry after a long hiatus period of personal reflection and research course re-evaluation and re-adjustment.
The present paper is lovingly dedicated to my parents, George and Helen Raptis, whose unceasing moral and material support of my research quests has never been diminished by the passage of time, no matter what its toll on both their ageing bodies and their lucid minds.
Last but not least, the unceasing ‘moral’ support of my lovely family: Kathleen, Francis, James and Cookie, is also warmly aknowledged, especially their patience and understanding in putting up with me over the years.